OpenStudy (anonymous):

Someone help me find the why and x intercept, 8x - 4y = 12

OpenStudy (saifoo.khan):

when x intercpts, y = 0 8x - 4y = 12 8x - 4(0) = 12 8x = 12 x = 12/8 x = 3/2

OpenStudy (saifoo.khan):

your name's like: C ME.

OpenStudy (saifoo.khan):

jk

OpenStudy (anonymous):

k last one lol, Rewrite this in slope intercept form? x – y = -3

OpenStudy (saifoo.khan):

post that on left.

OpenStudy (saifoo.khan):

Post 100, but on left.

OpenStudy (saifoo.khan):

we r here to help! :)